OpenStudy (anonymous):

2r - (5 - r) = 13 + 2r

OpenStudy (katrinakaif):

2r - (5 - r) = 13 + 2r 2r - 5 + r = 13 + 2r 3r - 5 = 13 + 2r 1r = 18 r = 18

OpenStudy (katrinakaif):

=) Tell me if you need additional help